# NAMET: Robust Massive Model Editing via Noise-Aware Memory Optimization

### v1

@@ B.2.1 CounterFact Metrics
    $-\frac{2}{3}\mathop{\sum}_{k}g_{2}(k)\log_{2}g_{2}(k)+\frac{4}{3}\mathop{\sum}_{k}g_{3}(k)\log_{2}g_{3}(k)$ (31)
@@
B.2.2 Wiki-recent Metrics
@@
    $-\frac{2}{3}\mathop{\sum}_{k}g_{2}(k)\log_{2}g_{2}(k)+\frac{4}{3}\mathop{\sum}_{k}g_{3}(k)\log_{2}g_{3}(k)$ (35)
@@ B.2.3 ZsRE Metrics
B.2.4 Examples of Evaluation Prompts with Prefixes
@@ B.3 Implementation Details
We implement all experiments on a single NVIDIA H800 GPU with 80GB memory. The details of our implementation across different models are outlined as follows:
  • LLAMA2-7B: We modify layers [3, 4] for editing factual knowledge. The hyperparameters $\lambda$s are set to [25000, 50000] respectively for two layers. During optimization, we iterate for 25 steps with 0.04 learning rate.
  • Qwen-2.5-7B: We modify layers [3, 4] for editing factual knowledge. The hyperparameters $\lambda$s are set to [500, 1000] respectively for two layers. During optimization, we iterate for 25 steps with 0.04 learning rate.
  • LLaMA2-13B: We modify layers [3, 4] for editing factual knowledge. The hyperparameters $\lambda$s are set to [25000, 50000] respectively for two layers. During optimization, we iterate for 25 steps with 0.04 learning rate.
  • Falcon-7B: We modify layers [3, 4] for editing factual knowledge. The hyperparameters $\lambda$s are set to [5000, 1000] respectively for two layers. During optimization, we iterate for 25 steps with 0.06 learning rate.
  • Deepseek-base-7B: We modify layers [3, 4] for editing factual knowledge. The hyperparameters $\lambda$s are set to [10000, 20000] respectively for two layers. During optimization, we iterate for 25 steps with 0.06 learning rate.
  • LLaMA3-8B: We modify layers [3, 4] for editing factual knowledge. The hyperparameters $\lambda$s are set to [7500, 15000] respectively for two layers. During optimization, we iterate for 25 steps with 0.008 learning rate.
Appendix C Algorithmic Description of NAMET
In this section, we provide a detailed algorithmic description of NAMET (Figure 4). Following previous work, we also spread MLP updates evenly across target critical layers $\mathcal{R}$. We define $L$ as the final layer of candidate layers, where new memories are fully represented. For each target edit $(s_{i},r_{i},o_{i})$, we first compute the target hidden representation $z_{i}$ to replace the original representation $h_{i}^{L}$ for modifying new factual associations. During the optimization procedure, the padding strategy is changed to left-padding to introduce additional noise for more isolated target representations. We proceed to separately modify MLP modules at every critical layer $l\in\mathcal{R}$ with evenly distributed change $\delta_{i}=z_{i}-h_{i}^{L}$ for every target edit $i$.
Optimizing target memory $z_{i}$. For each subject-relation pair $(s_{i},r_{i})$, we concatenate them into prompt $tp(s_{i},r_{i})$. We proceed to augment each prompt with additional $N_{FP}-1$ random prefixes of fixed length. The original prompt, together with its augmented companions, is tokenized using a left-padding tokenizer, introducing [unk] tokens to the beginning of the original prompt (as shown in Figure 4). These [unk] tokens carry no semantic meaning and provide additional noise during optimization. The target representation $z_{i}$ is then optimized via:
$z_{i}=h_{i}^{L}+\underset{\delta_{i}}{\arg\min}\frac{1}{N_{FP}}\sum_{j=1}^{N_{FP}}-\log G_{(h_{i}^{L}+=\delta_{i})}[o_{i}^{t}\mid f_{j}\oplus tp(s_{i}^{t},r_{i}^{t})]$ (39)
where $\{f_{j}\oplus tp(s_{i},r_{i})\}$ concatenates random prefixes $f_{j}$ with templated prompt $tp(s_{i},r_{i})$ for each target edit $i$. $G_{(h_{i}^{L}+=\delta_{i})}$ denotes that we add change $\delta_{i}$ to the hidden state $h_{i}^{L}$ during the model execution.
Spreading $\delta_{i}$ across candidate layers. Since the editing representation at earlier layers affects that of subsequent layers, we perform edits in ascending layer order. For each edit, we distribute the desired change $\delta_{i}$ proportionally across all candidate layers. Specifically, the residual to be added at layer $l\in\mathcal{R}$ is defined as:
$r_{i}^{l}=\frac{z^{i}_{t}-h^{L}_{i}}{L-l+1}$ (40)
where $h^{L}_{i}$ denotes the hidden state at the final layer $L$ before editing layer $l$.
In addition to $r_{i}^{l}$, we require the corresponding key $k_{i}^{l}$ to update the MLP module. For each subject $s_{i}$, the key is computed as:
$k_{i}^{l}=\frac{1}{N_{FP}}\sum_{j=1}^{N_{FP}}k(f_{j}\oplus s_{i}^{t}),\quad\text{where}\quad k(x)=\sigma(W_{in}^{l}\gamma(h^{l-1}_{i}(x)))$ (41)
Here, $f_{j}$ is the same random prefix concatenated with the subject $s_{i}$ mentioned above. Notably, prepending [unk] tokens does not affect key computation, thereby preserving the accuracy of the extracted key representations. After aggregating keys $K_{t}=[k_{1}^{t}\ |\ k_{2}^{t}\ |\ ...\ |\ k_{N_{t}}^{t}]$ and residuals $R_{t}=[r_{1}^{t}\ |\ r_{2}^{t}\ |\ ...\ |\ r_{N_{t}}^{t}]$ of all $N_{t}$ target edits, individual $\Delta^{l}$ for candidate layers could then be computed using Equation 4. We summarize the full algorithm in Algorithm C.1.
We provide an intuitive analysis of the effect of introducing [unk] tokens during optimization by examining the norm of hidden states at layer 4, as shown in Table 5. The results indicate that incorporating [unk] tokens introduces additional noise into the hidden states, increasing their norm by at least 4$\times$. Consequently, optimizing over these noisier representations can lead to more isolated and disentangled target representations.
Table 5: Hidden state norm of NAME and MEMIT across different models.
Model NAMET MEMIT
LLaMA2-7B 1321.3±114.6 6.7±0.9
Qwen-2.5-7B 21818.1±2399.9 45.7±5.7
LLaMA2-13B 670.5±26.5 7.3±0.9
Falcon-7B 257.4±2.9 32.0±3.4
Deepseek-7B 6275.6±192.9 23.9±2.8
LLaMA3-8B 570.2±25.4 3.7±0.6
Gemma-7B 70.9±6.1 15.6±2.8
Phi-1.5 209.5±14.2 48.3±5.6
@@ Appendix C Algorithmic Description of NAMET
2Data: Requested edits =$\mathcal{E}=\{(s_{i},r_{i},o_{i})\}$, generator $G$, layers to edit $\mathcal{S}$, covariances $C^{l}$
@@ 3 Result: Modified generator containing edits from $\mathcal{E}$
5 tokenizer.padding_side $\leftarrow$ "left"
6  // Change padding strategy to left
7 for $s_{i},r_{i},o_{i}\in\mathcal{E}$ do
8       optimize $\delta_{i}\leftarrow\arg\min_{\delta_{i}}\frac{1}{N_{FP}}\sum_{j=1}^{N_{FP}}-\log\mathbb{P}_{G_{(h_{i}^{L}+=\delta_{i})}}[o_{i}\mid f_{j}\oplus tp(s_{i},r_{i})]$
9       $z_{i}\leftarrow h_{i}^{L}+\delta_{i}$
@@
11
12for $l\in\mathcal{R}$ do
@@ 15             $k_{i}^{l}\leftarrow k_{i}^{l}=\frac{1}{N_{FP}}\sum_{j=1}^{N_{FP}}k(f_{j}\oplus s_{i})$
16             $r_{i}^{l}\leftarrow\frac{z_{i}-h_{i}^{L}}{L-l+1}$
17              // Distribute over remaining layers
18
@@ 20       $R^{l}\leftarrow[r_{1}^{l},\ldots,r_{N_{t}}^{l}]$
21       $\Delta^{l}\leftarrow R^{l}K^{l^{T}}(C^{l}+K^{l}K^{l^{T}})^{-1}$
22       $W^{l}\leftarrow W^{l}+\Delta^{l}$
23        // Update layer $l$ MLP weights in model
24
Algorithm C.1 The NAMET Algorithm
Appendix D Additional Experimental Results
In this section, we present additional experimental results and findings to further demonstrate the effectiveness of NAMET. First, we evaluate editing performance across different semantic categories. Next, we show that NAMET minimally interferes with the model’s general capabilities by assessing performance on the GLUE benchmarks. We then report editing performance on two additional LLMs: Gemma-7B[29] and Phi-1.5[16]. Finally, we provide a comprehensive analysis of the hyperparameters.
D.1 Editing Performance Involving Different Semantics
@@
Additional Finding 1. NAMET achieves superior editing performance across different semantics. We extract samples with specific relation types from the CounterFact dataset to evaluate the performance of different editing methods across semantic categories. As shown in Figure 7, NAMET consistently achieves the highest editing efficacy and generalization on both LLaMA2-7B and Qwen2.5 for most semantic types. On LLaMA2-7B, NAMET outperforms the second-best method (PMET) by approximately 10% in efficacy and 20% in generalization. On the “twin city” relation, NAMET achieves twice the efficacy and four times the generalization of PMET. MEMIT, in contrast, nearly fails to edit facts with this relation, with both efficacy and generalization scores close to 0%. In terms of editing specificity, NAMET performs better on Qwen2.5 than on LLaMA2. On LLaMA2, PMET surpasses NAMET by an average of 5%, whereas on Qwen2.5, NAMET achieves the highest specificity on 6 out of 8 relation types.
D.2 General Ability of Edited Models on GLUE Benchmarks
@@
Additional Finding 2. NAMET better preserves the general ability of LLMs after massive editing. We assess whether massive editing degrades the general capabilities of LLMs when using MEMIT, PMET, and NAMET. To this end, we evaluate the performance of three LLMs (LLaMA2-7B, Qwen2.5-7B, and LLaMA3-8B) on the GLUE benchmark after editing 10,000 knowledge facts. For reference, we also report the performance of each unedited (pre-edit) model. As shown in Figure 8, models edited with NAMET exhibit the smallest performance deviation from their pre-edited counterparts. On LLaMA3-8B, the average deviation across six GLUE metrics for the NAMET-edited model is 0.118, which is 0.04 lower than MEMIT and half that of PMET. On LLaMA2-7B, NAMET maintains deviations below 0.05 for 4 out of 6 metrics, with a maximum deviation of only 0.071. The overall average deviation is 0.036, which is the lowest among all methods. NAMET demonstrates a strong ability to preserve general capability on Qwen2.5-7B, with an average deviation of 0.189, significantly lower than those of MEMIT (0.269) and PMET (0.308).
We attribute this robustness to NAMET’s ability to extract more dispersed memory representations across knowledge items. Such dispersion reduces the likelihood of embedding collisions during massive editing, which may compromise the model’s general capabilities. By mitigating this interference, NAMET effectively preserves the original functionality of the LLM.
D.3 Analysis of Editing Performance across LLMs
@@
Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Flu.$\uparrow$ Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$
Gemma-7B MEMIT 93.01 54.33 74.88 5.38 84.61 73.66 23.04
PMET 91.69 47.24 75.22 5.33 83.05 73.95 23.85
NAMET 95.29 68.53 66.22 5.30 85.83 78.23 23.37
Phi-1.5 MEMIT 49.00 29.67 63.71 5.68 49.76 38.54 20.36
PMET 38.49 20.75 67.01 5.79 32.00 24.47 21.54
NAMET 67.76 40.13 62.08 5.80 51.54 41.02 21.29

### v4

@@ D.2.1 CounterFact Metrics
    $\frac{2}{3}\mathop{\sum}_{k}g_{2}(k)\log_{2}g_{2}(k)+\frac{4}{3}\mathop{\sum}_{k}g_{3}(k)\log_{2}g_{3}(k)$ (37)
@@
D.2.2 Wiki-recent Metrics
@@
    $\frac{2}{3}\mathop{\sum}_{k}g_{2}(k)\log_{2}g_{2}(k)+\frac{4}{3}\mathop{\sum}_{k}g_{3}(k)\log_{2}g_{3}(k)$ (41)
@@ D.2.3 ZsRE Metrics
D.2.4 Examples of Evaluation Prompts with Prefixes
@@ D.3 Implementation Details
We implement all experiments on a single NVIDIA H800 GPU with 80GB memory. During optimization, we iterate for 25 steps with 0.5 learning rate. We set $M=50$ for balancing the fine-grained alignment and optimization efficiency. The details of our implementation across different models are outlined as follows:
[…]
We justify our choice of editing layers by comparing against the common settings used in prior work (Meng et al., 2023; Li et al., 2025; Meng et al., 2022). On LLaMA2-7B, we evaluate EAMET with different layer combinations when editing 10,000 facts from CounterFact and ZsRE. As shown in Table˜7, EAMET achieves higher efficacy and generalization with layers [3, 4] compared to [4, 5, 6, 7, 8].
Table 7: Performance comparison of EAMET on LLAMA2-7B with different layer selections.
Layers Counterfact ZsRE
[…]
3, 4 89.09 61.21 72.19 519.23 89.47 81.34 15.70
4, 5, 6, 7, 8 77.58 36.83 73.43 516.63 87.14 76.91 15.92
This effect arises because, as the edited layer becomes deeper, the similarity between key embeddings of different knowledge items increases. As shown in Figure˜7, the average similarity across layers of LLaMA2-7B grows with layer depth. When the last edited layer is $8$, the average similarity is nearly twice that of layer $4$. This growth makes it more difficult to align the memory embedding space with the key embedding space, since KL divergence primarily captures distributional differences and we only apply MSE loss to the top-$M$ cosine similarities. Applying MSE to all cosine similarities may lead to vanishing gradients. As the number of similarities grows, the strongest ones become diluted, which slows convergence and hinders optimization. Such misaligned memory embeddings can substantially degrade both the effectiveness and robustness of massive editing.
Figure 7: Average of key similarities across different layers of LLaMA2-7B when editing 500 knowledge items.
Appendix E Algorithmic Description of EAMET
@@
2Data: Requested edits $\mathcal{E}=\{(s_{i},rel_{i},o_{i})\}$, generator $G$, layers to edit $\mathcal{S}$, covariances $C^{l}$
@@ 3 Result: Modified generator containing edits from $\mathcal{E}$
5 $K\leftarrow[]$
6 $L\leftarrow\text{final layer of candidate layers }\mathcal{R}$
7 for $s_{i},rel_{i},o_{i}\in\mathcal{E}$ do
9 $K\leftarrow K\cup\{k_{i}^{l}\}$
@@
[…]
12 $R\leftarrow[]$
13 for $s_{i},rel_{i},o_{i}\in\mathcal{E}$ do
 $r_{i}\leftarrow h_{i}^{L}$
 // Initialize $r_{i}$ as the original hidden state
14 $P_{r}^{(i)}\leftarrow\{P_{r}^{(i,j)}\mid j<i,r_{j}\in R\}$
15 $\bar{P}_{k}^{(i)}\leftarrow\{P_{k}^{(i,j)}\mid j<i,k_{j}\in K\}$
16 $L_{\text{KL}}(i)=\text{KL}\!\left(P_{r}^{(i)}\,\|\,\bar{P}_{k}^{(i)}\right)$
[…]
18 $L_{\text{MSE}}(i)=\frac{1}{M}\sum_{j\in I_{K}}\big\|P_{r}^{(i,j)}-P_{k}^{(i,j)}\big\|^{2}$
19 $r_{i}\leftarrow\arg\min_{r_{i}}\frac{1}{N_{FP}}\sum_{j=1}^{N_{FP}}-\log\mathbb{P}_{G_{(h_{i}^{L}+=r_{i})}}[o_{i}\mid f_{j}\oplus tp(s_{i},r_{i})]+\lambda_{KL}L_{\text{KL}}(i)+\lambda_{MSE}L_{\text{MSE}}(i)$
20 $z_{i}\leftarrow h_{i}^{L}+r_{i}$
21 $R\leftarrow R\cup\{r_{i}\}$
22
23
24for $l\in\mathcal{R}$ do
@@ 27    $k_{i}^{l}\leftarrow k_{i}^{l}=\frac{1}{N_{FP}}\sum_{j=1}^{N_{FP}}k(f_{j}\oplus s_{i})$
    $r_{i}^{l}\leftarrow\frac{z_{i}-h_{i}^{L}}{L-l+1}$
    // Distribute over remaining layers
28
@@ 30 $R^{l}\leftarrow[r_{1}^{l},\ldots,r_{N_{t}}^{l}]$
31 $\Delta\leftarrow R^{l}K^{l^{T}}(C^{l}+K^{l}K^{l^{T}})^{-1}$
 $W^{l}\leftarrow W^{l}+\Delta$
 // Update layer $l$ MLP weights in model
32
Algorithm 1 The EAMET Algorithm
In this section, we present a detailed description of the EAMET algorithm in Algorithm 1. The procedure consists of three main stages: (1) key embedding preparation, (2) aligning memory embeddings with key embeddings, and (3) distributing MLP updates across candidate layers.
Key Embedding Preparation. Following prior work (Meng et al., 2022; 2023; Li et al., 2025), we evenly distribute MLP updates across the critical target layers $\mathcal{R}$. We denote by $L$ the final candidate layer where new memories are fully represented (Line 5). Before optimizing residual embeddings, we first compute the key embeddings for each target edit (Line 7). To improve generalization, each subject is augmented with $N_{FP}-1$ random prefixes of fixed length $f_{i}$. All resulting key embeddings are aggregated into a matrix $K$ (Line 8). We then compute pairwise cosine similarities among key embeddings to obtain $P_{k}$ (Line 9).
Aligning Memory Embeddings with Key Embeddings. For each target edit, we initialize the residual embedding with the original hidden state (Line 12), since it naturally corresponds to the associated key embedding and thus provides a good starting point. We compute cosine similarities among residual embeddings to form $P_{r}^{(i)}$ (Line 13), and extract the corresponding key embedding structure $\bar{P}_{k}^{(i)}$ (Line 14). Alignment is achieved by minimizing the KL divergence between $P_{r}^{(i)}$ and $\bar{P}_{k}^{(i)}$ (Line 15). To further refine alignment, we select the indices $I_{K}$ corresponding to the top $M$ similarities in $P_{k}^{(i)}$ (Line 16) and minimize the MSE loss between $P_{r}^{(i)}$ and $P_{k}^{(i)}$ restricted to these indices (Line 17). The optimized residual embedding $r_{i}$ is obtained by minimizing this combined objective (Line 18) and stored in the set $R$ (Line 19).
Distributing MLP Updates Across Candidate Layers. We update MLP modules sequentially across layers $l\in\mathcal{R}$, as earlier edits affect subsequent representations. For each candidate layer, we compute key embeddings $k_{i}^{l}$ for all edits (Line 24) and residual embeddings $r_{i}^{l}$, distributing them proportionally across layers (Line 25). These embeddings are then aggregated into $K^{l}$ and $R^{l}$ (Lines 26-27) to update the MLP weights with $\Delta$ (Lines 28-29).
Appendix F Additional Experimental Results
In this section, we present additional experiments and findings to further validate the effectiveness of EAMET. We begin by evaluating editing performance across different semantic categories. Next, we assess its impact on the model’s general capabilities using the GLUE benchmarks. We then report results on two additional LLMs, Gemma-7B (Team et al., 2024) and Phi-1.5 (Li et al., 2023). We also examine how the order of edits affects EAMET’s performance. Furthermore, we explore its integration with sequential editing, showing that embedding alignment enables larger batch sizes per step and thus reduces the number of steps needed to edit the same set of knowledge items. Finally, we provide an ablation study on combining KL loss and MSE loss, along with a comprehensive analysis of the hyperparameters $\lambda_{KL}$, $\lambda_{MSE}$, and $M$.
F.1 Editing Performance Involving Different Semantics
@@
Additional Finding 1. EAMET Achieves Superior Editing Performance Across Different Semantics. We extract samples with specific relation types from the CounterFact dataset to evaluate the performance of different editing methods across semantic categories. As shown in Figure˜8, EAMET consistently achieves the highest editing efficacy and generalization on both LLaMA2-7B and Qwen2.5 for most semantic types. On LLaMA2-7B, EAMET outperforms the second-best method (PMET) by approximately 10% in efficacy and 20% in generalization. In terms of editing specificity, EAMET performs better on Qwen2.5 than on LLaMA2. On LLaMA2, PMET surpasses EAMET by an average of 5%, whereas on Qwen2.5, EAMET achieves the highest specificity on 6 out of 8 relation types.
[…]
F.2 General Ability of Edited Models on GLUE Benchmarks
@@
Additional Finding 2. EAMET Better Preserves the General Ability of LLMs After Massive Editing. We examine whether large-scale editing degrades the general capabilities of LLMs under MEMIT, PMET, and EAMET. Specifically, we evaluate three models (LLaMA2-7B, Qwen2.5-7B, and LLaMA3-8B) on the GLUE benchmark (Wang et al., 2018) after editing 10,000 knowledge facts from CounterFact and ZsRE. For reference, we also report the performance of the unedited models. As shown in Figure˜9, EAMET consistently yields the smallest performance deviation from pre-edit baselines. On Qwen2.5-7B, the average deviation across six GLUE tasks is only 0.083 for CounterFact and 0.032 for ZsRE, substantially lower than MEMIT (0.266 and 0.349) and PMET (0.310 and 0.276). A similar trend holds for LLaMA3-8B and LLaMA2-7B: on CounterFact, EAMET achieves average deviations of 0.083 and 0.025, compared to 0.155 and 0.043 for MEMIT, the second-best method.
We attribute this robustness to EAMET’s ability to extract more aligned memory representations across knowledge items. Such alignment reduces the likelihood of embedding inconsistency between key and residual spaces during massive editing, which may compromise the model’s general capabilities. By mitigating this interference, EAMET effectively preserves the original functionality of the LLM.
F.3 Evaluation on Additional LLMs
@@
[…]
Gemma-7B MEMIT 93.01 54.33 74.88 538.92 84.61 73.66 23.04
PMET 91.69 47.24 75.22 533.62 83.05 73.95 23.85
EAMET 95.29 68.53 70.22 530.93 91.69 86.43 23.37
Phi-1.5 MEMIT 49.00 29.67 63.71 568.21 49.76 38.54 20.36
PMET 38.49 20.75 67.01 579.19 32.00 24.47 21.54
EAMET 67.76 40.13 62.08 580.92 60.61 45.42 21.29